OpenStudy (anonymous):

What is the general form of the equation for the given circle centered at o (0,0) (4,5) x^2+y^2+x+y-41=0

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[(x-0)^2 +(y-0)^2=16+25=41\\ x^2+y^2 -41=0 \]

OpenStudy (anonymous):

x^2+y^2-41=0

OpenStudy (anonymous):

yes